OpenStudy (anonymous):

19. Suppose a and b vary inversely, and b = 8 when a = 6. Write a function that models the variation and find b when a = 30.

terenzreignz (terenzreignz):

Hello :) When two values vary inversely, it means their product must remain constant :) b = 8 when x = 6 , right? ab = 6 x 8 = 48 When a = 30, we want to find b... but remember, whatever that b is, when multiplied to 30, it must STILL be 48, so... 30b = 48 Now find b :)

OpenStudy (anonymous):

to find b would i divide 48 by 30????

terenzreignz (terenzreignz):

That's right :)

OpenStudy (anonymous):

1.6=b is that the right answer?

terenzreignz (terenzreignz):

1.6 x 30 = 48... yep, it's correct :)

OpenStudy (anonymous):

thank you can you help me with a few more questions please?

terenzreignz (terenzreignz):

Go ahead.

OpenStudy (anonymous):

6. Find any points of discontinuity for the rational function. y = x-8/x^2 +6x-7 (1 point) x = 1, x = 7 x = 8 x = 1, x = –7 x = –1, x = 7 7. Describe the vertical asymptote(s) and hole(s) for the graph of y = x-5/x^2 +4x +3 asymptotes: x = –3, –1 and no holes. asymptote: x = –3 and hole: x = –5 asymptotes: x = –3, –1 and hole: x = –5 asymptote: x = –5 and hole: x = –3 8. Find the horizontal asymptote of the graph of y = -4x^6 +6x +3/8x^6 +9x+3 (1 point) y = 1 y = –1/2 y = 0 There is no horizontal asymptote.

terenzreignz (terenzreignz):

For number 6, it's a rational function, which is continuous everywhere on its domain, so your best bet is to find values of x which are not in its domain, namely, where the denominator is zero.